OpenStudy (anonymous):

For 2x – 3y = –7, when x = 4, What is the value of y?

OpenStudy (anonymous):

13/3

OpenStudy (phoenixfire):

Rearrange the equation for y. Then substitute 4 for x.

OpenStudy (anonymous):

-3y = -7 - 6 -> -3y=-13 -> to cancel sa 3. you should divide it both side. and you will get the answer.

OpenStudy (phoenixfire):

\[2x-3y=-7\]\[2x+7=3y\]\[2(4)+7=3y\]\[15=3y\]\[{15 \over 3}=y\]\[y=5\]

OpenStudy (anonymous):

HAHHHAAHHHAA. i'm so sorry i was confused it should be 8 not 6. xD